# Jet schemes, Newton polygons and continued fractions

### v2

@@ 6. Useful Results
The inclusion: $F^{\alpha_{i}}_{i,k_{i}}\cap\mathbf{A}^{2}_{m}\subseteq\overline{F^{\alpha_{j}}_{j,k_{j}}\cap\mathbf{A}^{2}_{m}}$ tells us that
$1+m-\nu_{\alpha_{i}}(f)+|\alpha_{i}|\geq 1+m-\nu_{\alpha_{j}}(f)+|\alpha_{j}|$
by comparing codimensions. Then, we get the inequality by simplifying:
$|\alpha_{i}-\alpha_{j}|=|\alpha_{i}|-|\alpha_{j}|\geq\nu_{\alpha_{i}}(f)-\nu_{\alpha_{j}}(f)$ (3)
@@
As $U:=\mathbf{A}^{2}_{m}\cap F^{\alpha_{i}}_{i,k_{i}}\subseteq\mathcal{H}^{\alpha_{j}}_{m}$ and $x^{(\alpha_{i,1})}\neq 0$ in $U$ then $\alpha_{i,1}\geq\alpha_{j,1}$ and similarly $\alpha_{i,2}\geq\alpha_{j,2}$ using $y^{(\alpha_{i,2})}\neq 0$ in $U$. Choose $\eta\in supp(f)$ such that $\eta\cdot\alpha_{i}=\nu_{\alpha_{i}}(f)$ and set $\delta=\alpha_{i}-\alpha_{j}\in\mathbf{N}^{2}_{0}$. Since $\eta\in supp(f)$ then $\eta\cdot\alpha_{j}\geq\nu_{\alpha_{j}}(f)$.
Then,
$|\delta|=|\alpha_{i}-\alpha_{j}|\geq\nu_{\alpha_{i}}(f)-\nu_{\alpha_{j}}(f)\geq\eta\cdot\alpha_{i}-\eta\cdot\alpha_{j}=\eta\cdot\delta$
So, $|\delta|\geq\nu_{\delta}(f)$. Choose $(a,b)\in supp(f)$ such that $\nu_{\delta}(f)=a\delta_{1}+b\delta_{2}\leq\delta_{1}+\delta_{2}$. This implies that $a=0$ or $b=0$ but not both, since $f(0)=0$. Since $a$ and $b$ play symmetric roles then suppose $a=0$ with $b\neq 0$. So, $b\delta_{2}\leq\delta_{1}+\delta_{2}$. Consider two cases:
  • If $\delta_{2}>\delta_{1}$ then $2\delta_{2}>\delta_{1}+\delta_{2}\geq b\delta_{2}$ thus $2>b$ and $b=1$. Then $(0,1)\in supp(f)$. Since we can take $f$ to be Weierstrass with respect to $y$, then, $f=y+a(x)$. $\frac{\partial f}{\partial y}=1$, then $(f=0)$ is not singular at the origin. Hence the contradiction.
  • If $0<\delta_{2}\leq\delta_{1}$, let’s get back to $|\delta|\geq\eta\cdot\delta$. Thus, $\delta_{1}(1-\eta_{1})+\delta_{2}(1-\eta_{2})\geq 0$.
    $(2-\eta_{1}-\eta_{2})\delta_{1}\geq 0$
    Then $\eta_{1}+\eta_{2}=2$ then $\eta_{1}=\eta_{2}=1$ since both are $\geq 1$. Since $\nu_{\alpha_{i}}(f)=\eta\cdot\alpha_{i}$, this shows $\nu_{\alpha_{i}}(f)=|\alpha_{i}|$. Then, $\alpha_{i}\cdot[\eta-(1,1)]=0$. Then, $\alpha_{i}\perp\eta-(1,1)$. Moreover, if $\eta_{1}=0$ or $\eta_{2}=0$ then $in_{\alpha_{i}}(f)$ is a monomial, which is absurd with the fact that $\alpha_{i}\in T(f)$. So, $\eta\geq(1,1)$. Now, as $\alpha_{i}$ and $\eta-(1,1)$ are both in the same quadrant $\mathbf{R}^{2}_{\geq 0}$ they can’t be orthogonal. Hence the contradiction.
  • Finally if $\delta_{1}\geq\delta_{2}=0$. Using $|\delta|\geq\eta\cdot\delta$, then again:
    $(2-\eta_{1}-\eta_{2})\delta_{1}\geq 0$
    If $\delta_{1}>0$ then $\eta_{1}=\eta_{2}=1$ since both are $\geq 1$. We find in the same way as the previous point a contradiction. So, $\delta_{1}=0$ and thus, $\delta_{1}=\delta_{2}=0$.
We found $\delta=(0,0)$. So, $\alpha_{i}=\alpha_{j}$. Now computing codimensions of the inclusion:
@@
we find that they have the same codimension (since $\alpha_{i}=\alpha_{j}$). Since the parts of the inclusion are irreducible with same codimension then they are equal. This completes the proof. ∎
Below, in the next lemmas and theorem, we show that $\mathcal{H}^{\gamma}_{m}$ and $\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha}_{i,k}}$ are not comparable by inclusion, defining then irreducible components of $\mathcal{C}^{0}_{m}$.
@@
$T(f_{i})=\mathbf{R}_{\geq 0}\alpha_{i}:\alpha_{i}\in\mathbf{N}^{2}$
@@
$\mathcal{H}^{\gamma}_{m}\supseteq\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha_{i}}_{i,k_{i}}}$
@@
As $\mathbf{A}^{2}_{m}\cap F^{\alpha_{i}}_{i,k_{i}}\subseteq\mathcal{H}^{\gamma}_{m}$ then $\alpha_{i}\geq\gamma$ for the product order. Then,
@@
Hence the contradiction. This concludes the proof. ∎
@@
$T(f_{i})=\mathbf{R}_{\geq 0}\alpha_{i}:\alpha_{i}\in\mathbf{N}^{2}$
@@
$\mathcal{H}^{\gamma}_{m}\subseteq\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha_{i}}_{i,k_{i}}}$
@@
When, $m=\nu_{\alpha_{i}}(f)$, this comes from the fact that
@@
is irreducible with codimension $|\alpha_{i}|+1$, the same as $\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha_{i}}_{i,k_{i}}}$. Now, suppose $m>\nu_{\alpha_{i}}(f)$ and $\mathcal{H}^{\gamma}_{m}\subsetneq\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha_{i}}_{i,k_{i}+h}}$. Then, by comparing codimensions:
$|\alpha_{i}|+2>|\alpha_{i}|+1+h$
Thus, $1>h=m-\nu_{\alpha_{i}}(f)>0$. Hence the contradiction. ∎
@@
$T(f_{i})=\mathbf{R}_{\geq 0}\alpha:\alpha\in\mathbf{N}^{2}$
and $\nu_{\alpha}(f)\leq m$. Set $k=\nu_{\alpha}(f_{i})+m-\nu_{\alpha}(f)$. Let $\gamma\in\mathbf{N}^{2}$ be minimal for the product order, such that $\nu_{\gamma}(f)>m$ and $\gamma\neq\alpha+(1,1)$. Then, we have
$\mathcal{H}^{\gamma}_{m}\not\subset\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha}_{i,k}}$
@@
Suppose (to find a contradiction) the inclusion holds. Then $\gamma\geq\alpha$ component-wise. Let’s show $|\gamma|\leq 1+|\alpha|+m-\nu_{\alpha}(f)$ by induction on $m\geq\nu_{\alpha}(f)$.
  • if $m=\nu_{\alpha}(f)$ then by the minimality of $\gamma\geq\alpha$ we have either $\gamma=\alpha+(1,0)$ or $\gamma=\alpha+(0,1)$ because if we set $n_{i}=\nu_{\alpha}(f_{i})$, $r=\nu_{\alpha}(R)$ and $f=Rf_{i}$
@@
    with $Rad\langle R^{(r)}\rangle=\langle x^{(\alpha_{1})}y^{(\alpha_{2})}\cdot*\rangle$ where * is a product of pairwise distinct binomials. So this shows $|\gamma|=|\alpha|+1$.
@@
    $|\gamma|\leq 1+|\alpha|+m+1-\nu_{\alpha}(f)$
  • if $m>\nu_{\alpha}(f)$ and $\nu_{\gamma}(f)=m+1$ then $\gamma$ (which is minimal) either bifurcates to $\gamma+(1,0)$ or $\gamma+(0,1)$ or gives rise to a component when $\gamma\in T(f)$
    $\overline{\mathbf{A}^{2}_{m}\cap F^{\gamma}_{j,\nu_{\gamma}(f_{j})+1}}$
    whose codimension is $|\gamma|+1$. If we are in the first or second case, since $\nu_{\gamma}(f)<\nu_{\gamma+\epsilon}(f)$ then at level $m+2$ we have:
@@
    with $\gamma+\epsilon$ minimal. The last case when $\gamma\in T(f)$ tell us that
    $\mathbf{A}^{2}_{m}\cap F^{\gamma}_{j,\nu_{\gamma}(f_{j})+1}\subset\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha}_{i,k}}$
    which contradicts the previous theorem.
So, we have shown $|\gamma|\leq 1+|\alpha|+m-\nu_{\alpha}(f)$. Together with the inclusion which we supposed we find a clear contradiction (unless $\gamma=\alpha+(1,1)$). The case $\gamma=\alpha+(1,1)$ is excluded by hypothesis. ∎

### v3

@@ 6. Useful Results
The inclusion: $F^{\alpha_{i}}_{i,k_{i}}\cap\mathbf{A}^{2}_{m}\subseteq\overline{F^{\alpha_{j}}_{j,k_{j}}\cap\mathbf{A}^{2}_{m}}$ tells us (via comparing codimensions) that
$1+m-\nu_{\alpha_{i}}(f)+|\alpha_{i}|\geq 1+m-\nu_{\alpha_{j}}(f)+|\alpha_{j}|,$
which in turn leads to the following inequality
$|\alpha_{i}-\alpha_{j}|=|\alpha_{i}|-|\alpha_{j}|\geq\nu_{\alpha_{i}}(f)-\nu_{\alpha_{j}}(f).$ (3)
@@
Since $U:=\mathbf{A}^{2}_{m}\cap F^{\alpha_{i}}_{i,k_{i}}\subseteq\mathcal{H}^{\alpha_{j}}_{m}$ and $x^{(\alpha_{i,1})}\neq 0$ in $U$, then $\alpha_{i,1}\geq\alpha_{j,1}$ and similarly we conclude that $\alpha_{i,2}\geq\alpha_{j,2}$ as $y^{(\alpha_{i,2})}\neq 0$ in $U$. Now choosing $\eta=(a,b)\in supp(f)$ a vertex of the Newton polygon of $f$ such that $\eta\cdot\alpha_{i}=\nu_{\alpha_{i}}(f)$ and setting $\delta=\alpha_{i}-\alpha_{j}\in\mathbf{N}^{2}_{0}$, then $\eta\cdot\alpha_{j}\geq\nu_{\alpha_{j}}(f)$. Therefore,
$|\delta|=|\alpha_{i}-\alpha_{j}|\geq\nu_{\alpha_{i}}(f)-\nu_{\alpha_{j}}(f)\geq\eta\cdot\alpha_{i}-\eta\cdot\alpha_{j}=\eta\cdot\delta,$
and hence $a\delta_{1}+b\delta_{2}\leq\delta_{1}+\delta_{2}$. But if $a,b\geq 1$ then $\delta\cdot(a-1,b-1)=0$ where both $\delta$ and $(a-1,b-1)\in\mathbf{R}_{\geq 0}^{2}$ are in the first quadrant, which is obviously a contradiction. So $a=0$ or $b=0$ but not both, since $f(0)=0$. As $a$ and $b$ play symmetric roles, then we may suppose $a=0$ with $b\neq 0$ so that $b\delta_{2}\leq\delta_{1}+\delta_{2}$. As $f$ defines a singularity then $b\geq 2$, since otherwise $f=y+a(x)$ can be written in Weierstrass form defining a smooth curve. Therefore, $\delta_{1}\geq\delta_{2}(b-1)\geq\delta_{2}$. Now, let’s set $a=\nu_{\alpha_{i}}(f_{i})+m-\nu_{\alpha_{i}}(f)$ and for simplicity, $\alpha=\alpha_{i}$ and $\beta=\alpha_{j}$. We have
$\varnothing\neq F^{\alpha_{i}}_{i,a}\cap\mathbf{A}^{2}_{m}\subseteq\overline{F^{\alpha_{j}}_{j,k_{j}}\cap\mathbf{A}^{2}_{m}}\subseteq\mathcal{H}^{\alpha_{j}}_{m}\cap\{f^{[\nu_{\alpha_{j}}(f_{j})]}_{j}=0\},$
$(x^{(0)},...,x^{(\beta_{1}-1)},y^{(0)},...,y^{(\beta_{2}-1)},f_{j}^{[\nu_{\beta}(f_{j})]})\subseteq(I:x^{(\alpha_{1})}).$
where $I=(x^{(0)},...,x^{(\alpha_{1}-1)},y^{(0)},...,y^{(\alpha_{2}-1)},f_{i}^{[\nu_{\alpha}(f_{i})]},...,f^{(a)}_{i})\subseteq\mathcal{O}_{\mathbf{A}^{2}_{m}}$ and
$(I:x^{(\alpha_{1})})=\{g\in\mathcal{O}_{\mathbf{A}^{2}_{m}}:x^{(\alpha_{1})}g\in I\}$
is the colon ideal. We can write
$x^{(\alpha_{1})}f^{[\nu_{\beta}(f_{j})]}_{j}=\sum_{j=0}^{\alpha_{1}-1}x_{j}X_{j}+\sum_{j=0}^{\alpha_{2}-1}y_{j}Y_{j}+\sum_{j=\nu_{\alpha}(f_{i})}^{a}f^{(j)}_{i}\Phi_{j}.$
Let’s annihilate the variables,
$x^{(0)},...,x^{(\beta_{1}-1)},y^{(0)},...,y^{(\beta_{2}-1)},x^{(\alpha_{1}+1)},...,x^{(m)},y^{(\alpha_{2}+1)},...,y^{(m)},$
we obtain a power series ring with variables $x^{(\beta_{1})},...,x^{(\alpha_{1})},y^{(\beta_{2})},...,y^{(\alpha_{2})}$ and the expressions
$f_{i}^{[\nu_{\alpha}(f_{i})]}\equiv{[y^{(\alpha_{2})}]}^{p_{i}}-c_{i}{[x^{(\alpha_{1})}]}^{q_{i}},$
$f_{j}^{[\nu_{\beta}(f_{j})]}\equiv{[y^{(\beta_{2})}]}^{p_{j}}-c_{j}{[x^{(\beta_{1})}]}^{q_{j}}$
where $T(f_{l})=\mathbf{R}_{\geq 0}(p_{l},q_{l})$ and $gcd(p_{l},q_{l})=1$ for $l\in\{i,j\}$. Now, if $\delta_{1}>0$ i.e. $\alpha_{1}>\beta_{1}$, we annihilate $x^{(\beta_{1})},...,x^{(\alpha_{1}-1)}$ to obtain in a power series ring
$x^{(\alpha_{1})}[y^{(\beta_{2})}]^{p_{j}}=\sum_{j=\beta_{2}}^{\alpha_{2}-1}y_{j}Y_{j}+([{y^{(\alpha_{2})}]}^{p_{i}}-c_{i}{[x^{(\alpha_{1})}}^{q_{i}}]]\Phi.$
Next, we distinguish the following 2 cases:
  • if $\alpha_{2}=\beta_{2}$ then inside a power series we have an expression of the form
    $x^{(\alpha_{1})}[{y^{(\beta_{2})}}]^{p_{j}}=([{y^{(\alpha_{2})}]}^{p_{i}}-c_{i}{[x^{(\alpha_{1})}}^{q_{i}}])\Phi,$
    which is clearly not possible.
  • if $\alpha_{2}>\beta_{2}$, then annihilate the variables $y^{(\beta_{2})},...,y^{(\alpha_{2}-1)}$ so that in a power series ring, we have an expression of the form
    $x^{(\alpha_{1})}[{y^{(\beta_{2})}}]^{p_{j}}=([{y^{(\alpha_{2})}]}^{p_{i}}-c_{i}{[x^{(\alpha_{1})}}^{q_{i}}])\Psi.$
Both cases lead to a contradiction, so $0\leq\delta_{2}\leq\delta_{1}=0$. As $\delta=(0,0)$, then $\alpha_{i}=\alpha_{j}$. Now inspecting codimensions in the inclusion:
@@
shows that both have the same codimension (as $\alpha_{i}=\alpha_{j}$). Since each side of the inclusion is irreducible and both have the same codimension, then they must be equal. This completes the proof. ∎
Our next results deal with showing that $\mathcal{H}^{\gamma}_{m}$ and $\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha}_{i,k}}$ are not comparable by inclusion.
@@
$T(f_{i})=\mathbf{R}_{\geq 0}.\alpha_{i}:\alpha_{i}\in\mathbf{N}^{2}$
@@
$\mathcal{H}^{\gamma}_{m}\supseteq\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha_{i}}_{i,k_{i}}}.$
@@
If $\mathbf{A}^{2}_{m}\cap F^{\alpha_{i}}_{i,k_{i}}\subseteq\mathcal{H}^{\gamma}_{m}$ then $\alpha_{i}\geq\gamma$ for the product order. Therefore,
@@
which is a contradiction. This concludes the proof. ∎
@@
Let $f\in\mathcal{O}_{k^{2}}$ define a Newton plane curve. Let $f_{i}$ define a branch of $f$ such that
$T(f_{i})=\mathbf{R}_{\geq 0}.\alpha_{i}:\alpha_{i}\in\mathbf{N}^{2}$
@@
$\mathcal{H}^{\gamma}_{m}\subseteq\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha_{i}}_{i,k_{i}}}.$
@@
If $m=\nu_{\alpha_{i}}(f)$, then we can write
@@
as both sides are irreducible with the same codimension $|\alpha_{i}|+1$. For the second part, suppose that $m>\nu_{\alpha_{i}}(f)$ and $\mathcal{H}^{\gamma}_{m}\subsetneq\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha_{i}}_{i,k_{i}+h}}$. Then, by comparing codimensions:
$|\alpha_{i}|+2>|\alpha_{i}|+1+h,$
we arrive at the inequality $1>h=m-\nu_{\alpha_{i}}(f)>0$ which is a contradiction. ∎
@@
Let $f\in\mathcal{O}_{k^{2}}$ define a Newton plane curve. Let $f_{i}$ define a branch of $f$ such that
$T(f_{i})=\mathbf{R}_{\geq 0}.\alpha:\alpha\in\mathbf{N}^{2}$
and $\nu_{\alpha}(f)\leq m$. Set $k=\nu_{\alpha}(f_{i})+m-\nu_{\alpha}(f)$. Let $\gamma\in\mathbf{N}^{2}$ be minimal for the product order, such that $\nu_{\gamma}(f)>m$. Moreover, assume that, $\gamma\neq\alpha+(1,1)$. Then, it holds that
$\mathcal{H}^{\gamma}_{m}\not\subset\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha}_{i,k}}.$
@@
Suppose by contradiction that the inclusion holds. If the inclusion is an equality then $\varnothing\neq F^{\alpha}_{i,k}\cap\mathbf{A}^{2}_{m}\subset\mathcal{H}^{\gamma}_{m}$. We then have $x^{(\alpha_{1})}\neq 0$ and $y^{(\alpha_{2})}\neq 0$ in $\mathcal{H}^{\gamma}_{m}$ and $\gamma_{1}\leq\alpha_{1}$ and $\gamma_{2}\leq\alpha_{2}$. Since $\mathcal{H}^{\gamma}_{m}\subseteq\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha}_{i,k}}\subseteq\mathcal{H}^{\alpha}_{m}$, then $\gamma\geq\alpha$. But $m\geq\nu_{\alpha}(f)=\nu_{\gamma}(f)>m$. This is a contradiction. So the inclusion which we assumed has to be strict. Next, we show that $|\gamma|\leq 1+|\alpha|+m-\nu_{\alpha}(f)$ by induction on $m\geq\nu_{\alpha}(f)$:
  • At the level $m=\nu_{\alpha}(f)$ (since $\gamma\geq\alpha$ is minimal), $\alpha$ is deleted and replaced by $\gamma=\alpha+(1,1)$, $\gamma=\alpha+(1,0)$ or $\gamma=\alpha+(0,1)$ because if we set $n_{i}=\nu_{\alpha}(f_{i})$, $r=\nu_{\alpha}(R)$ and $f=Rf_{i}$, then
@@
    with $Rad\langle R^{(r)}\rangle\subset\langle x^{(\alpha_{1})}y^{(\alpha_{2})}\cdot*\rangle$ where * is a product of pairwise distinct binomials. But since $\gamma\neq\alpha+(1,1)$ then this proves that $|\gamma|=|\alpha|+1$.
@@
    $|\gamma|\leq 1+|\alpha|+m-\nu_{\alpha}(f)<1+|\alpha|+m+1-\nu_{\alpha}(f).$
  • If $m>\nu_{\alpha}(f)$ and $\nu_{\gamma}(f)=m+1$ then $\gamma$ (which is minimal) either bifurcates to one of these $\{\gamma+(1,0),\ \gamma+(0,1)\}$ or when $\gamma\in T(f)$ it gives rise to a component
    $\overline{\mathbf{A}^{2}_{m+1}\cap F^{\gamma}_{j,\nu_{\gamma}(f_{j})}}$
    whose codimension is $|\gamma|+1$. For the first case, as $\nu_{\gamma}(f)<\nu_{\gamma+\epsilon}(f)$ then at level $m+2$ we can write
@@
    where $\gamma+\epsilon$ is minimal. Let us denote $\pi:=\pi_{m+1,m}:\mathbf{A}^{2}_{m+1}\rightarrow\mathbf{A}^{2}_{m}$ the projection morphism. The second case when $\gamma\in T(f)$ tells us that
    $\mathbf{A}^{2}_{m+1}\cap F^{\gamma}_{j,\nu_{\gamma}(f_{j})}=\{f_{j}^{[\nu_{\gamma}(f_{j})]}=0\}\cap\pi^{-1}[\mathbf{A}^{2}_{m}\cap F^{\gamma}_{j,\nu_{\gamma}(f_{j})-1}]=\{f^{[\nu_{\gamma}(f_{j})]}_{j}=0\}\cap\pi^{-1}(\mathcal{H}^{\gamma}_{m}).$
    Since the inclusion $\mathcal{H}^{\gamma}_{m}\subsetneq\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha}_{i,k}}$ is strict, then
    $\{f^{[\nu_{\gamma}(f_{j})]}_{j}=0\}\cap\pi^{-1}(\mathcal{H}^{\gamma}_{m})\subsetneq\{f^{[\nu_{\gamma}(f_{j})]}_{j}=0\}\cap\overline{\mathbf{A}^{2}_{m+1}\cap F^{\alpha}_{i,k}},$
    $\mathbf{A}^{2}_{m+1}\cap F^{\gamma}_{j,\nu_{\gamma}(f_{j})}\subsetneq\overline{\mathbf{A}^{2}_{m+1}\cap F^{\alpha}_{i,k}}\cap\{f^{[\nu_{\gamma}(f_{j})]}_{j}=0\},$
    $\overline{\mathbf{A}^{2}_{m+1}\cap F^{\gamma}_{j,\nu_{\gamma}(f_{j})}}\subsetneq\mathcal{H}^{\alpha}_{m+1}\cap\{f_{i}^{[\nu_{\alpha}(f_{i})]}=...=f_{i}^{(k)}=f^{[\nu_{\gamma}(f_{j})]}_{j}=0\}.$
    As the inclusion is strict, then by comparing codimensions of the irreducible sets, we obtain
    $1+|\gamma|>2+|\alpha|+m-\nu_{\alpha}(f),$
    which contradicts the induction hypothesis: $|\gamma|\leq 1+|\alpha|+m-\nu_{\alpha}(f)$. Therefore, the case $\gamma\in T(f)$ can not occur.
Thus, we have shown $|\gamma|\leq 1+|\alpha|+m-\nu_{\alpha}(f)$, and the proof is complete.